OpenStudy (anonymous):

Describe the domain and range of the linear transformation defined by the matrix A=(1 3; 1 1; 2 3) What is the rank of matrix A?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Is it because there are two rows in the matrix?

OpenStudy (anonymous):

rank = min(m,n) if you had \[\left[\begin{matrix}1 & 3 & 1 \\ 1 & 2 &3\end{matrix}\right]\] what will you get?

OpenStudy (anonymous):

I've actually never learned rank, I'm trying to help friends study for test and it's covering some material I haven't done :/

OpenStudy (anonymous):

I have to run to a meeting, but I'll be back on in an hour or so. Thanks again, I appreciate all the help :)

OpenStudy (anonymous):

rank is the min (row, column) of the nonzero rows \[\left[\begin{matrix}1 & 1 \\ 0 & 0\end{matrix}\right]\rightarrow rank = 1 \] you had 2rows and two columns put a row full of zeros so it will be 1 row and 2 columns; the min is 1 \[\left[\begin{matrix}1 & 0 & 3 & 4\\ 0 & 0 & 5 & 6 \\ 1 & 0 & 9 & 7 \\0 & 0 &1 & 0 \end{matrix}\right]\rightarrow rank =3 \] you had 4rows and 4 columns put a column full of zeros so it will be 4 row and 3 columns; the min is 3 do you understand it now?

OpenStudy (anonymous):

I think so, thanks :)